# 3D Visual Illusion Depth Estimation

### v2

@@ 4.1 Dual-Branch Prediction Network
As shown in Figure 4, our dual-branch prediction network comprises a binocular disparity estimation branch and a monocular depth estimation branch. The stereo branch adopts an iterative optimization framework, extracting features from rectified image pairs and constructing a cost volume via dot product. Starting from an initial disparity of zero, a GRU-based module iteratively refines the disparity map. The monocular branch takes the left image as input and uses the fixed-weight DepthAnything V2 [43] to predict affine-invariant inverse depth, representing normalized relative disparity. We also extract features from the penultimate layer of DepthAnything V2 and adapt them via learnable convolutional layers. These adapted features serve as left-view context to guide disparity refinement in the stereo branch.
@@ 4.2 VLM-Based Fusion Network
VLM Prediction Stage  In this stage, we utilize the pretrained QwenVL2-7B model [37, 1]. The visual prompts consist of the left-view RGB image, binocular disparity map, and monocular disparity map. Due to the complexity and variability of monocular texture-based ambiguities, which are hard to describe precisely in text, the textual prompts are designed primarily from a binocular perspective. Specifically, they focus on reflective, transparent, or other material types that typically cause stereo matching failures. By leveraging the vision-language model’s general reasoning capabilities, we extract embedding features that help distinguish between monocular and binocular depth cues under language guidance.
@@
Specifically, the textual prompts are first mapped into the embedding space. The image and text embeddings are concatenated to form $c_{e}$, which is combined with the intermediate state $y_{t_{k}}$ and passed through a stack of Transformers. Cross-attention is used to inject conditional information, and the velocity field $v_{t_{k}}$ is predicted. After multiple iterations via Equation 8, the final state $y_{1}$ is tensorized into a 2D format and decoded into image space using a variational autoencoder. The resulting feature is then concatenated with the cost volume and passed through convolutional layers to predict the confidence map $I_{c}$:
@@
where $\sigma$ is the sigmoid function, $\mathcal{F}_{c}$ denotes convolution, $\mathcal{G}_{c}$ is the VAE decoder, $y^{\prime}_{1}$ is the tensorized form of $y_{1}$, and $V(D_{s})$ is the cost volume sampled around binocualr disparity $D_{s}$.
Global Fusion Stage  Gloabal fusion first aligns monocualr disparity $D_{M}$ to the absolute disparity space and then fuse the aligned monocualr disparity $\tilde{D}_{m}$ with binocualr disaprity $D_{s}$. The alignment is achieved by affine transformation parameters $s_{m},t_{m}$:
@@
$s_{m}$ and $t_{m}$ are learned through convolutions on the concatenation of the monocular disparity $D_{m}$ and the binocular disparity $D_{s}$. Since $s_{m}$ and $t_{m}$ on low-confident regions is unrealible, we further refine the parameters by pooling $s_{m}$ and $t_{m}$ of the high-confident neighbors. After acquiring refined parameters $s_{m}$ and $t_{m}$, we compute the aligned monocular disparity $\tilde{D}_{m}$ using Equation 10. $\tilde{D}_{m}$, $D_{s}$, and $I_{c}$ are then concatenated and passed through convolutions and upsampling layers to generate the final high-resolution disparity map.
@@ 5 Experiment
We first train the models on the SceneFlow dataset [27] and then fine-tune them on our 3D-Visual-Illusion dataset. The finetuned models are then tested on the 3D-Visual-Illusion test set and the Booster training set [30]. We compare with monocular depth estimation methods (DepthAnything V2 [43], Marigold [17], Metric3D [44], and DepthPro [3]), multi-view large model methods (Dust3R [38] and VGGT [36]), and stereo matching methods (RAFT-Stereo [26], Selective-IGEV [39], and MochaStereo [5]). For more training details, loss, evaluation metrics, and prompts, please refer to the supplementary material.
@@
Table 1: Evaluation comparison in disparity and depth space on the 3D-Visual-Illusion dataset. ‘+ align’ indicates alignment using globally shared affine parameters computed from ground truth.
Method Disparity Space Depth Space
EPE $\downarrow$ bad2 $\downarrow$ bad3 $\downarrow$ bad5 $\downarrow$ AbsRel $\downarrow$ RMSE $\downarrow$ $\delta_{1}$uparrow
DepthAnything V2 [43] 5.81 61.45 43.18 30.57 0.14 0.15 92.86
Metric3D [44] 12.46 94.11 91.14 82.05 0.34 0.29 48.97
DepthAnything V2 metric [43] 16.24 92.53 87.43 75.25 0.52 0.39 48.75
DepthPro [3] 12.26 87.08 80.60 62.43 0.28 0.25 65.92
Marigold [17] 21.16 65.67 59.67 53.19 0.45 0.37 63.65
DepthAnything V2 metric [43] + align 5.23 56.82 45.50 28.89 0.17 0.15 93.70
Metric3D [44] + align 5.70 66.26 50.92 40.43 0.17 0.17 94.80
DepthPro [3] + align 4.36 44.98 34.98 24.70 0.09 0.10 93.83
Dust3R [38] 6.74 52.89 45.31 36.61 0.25 0.22 87.09
VGGT [36] 6.16 53.32 44.89 37.20 0.13 0.12 78.46
RAFT-Stereo [26] 1.62 24.32 13.20 2.97 0.04 0.06 99.18
Selective-RAFT [39] 1.58 23.46 12.65 2.57 0.03 0.07 99.60
Selective-IGEV [39] 1.67 24.06 13.11 2.99 0.04 0.10 99.26
MochaStereo [5] 1.75 25.49 14.11 3.54 0.04 0.11 98.76
StereoAnything [13] 2.41 29.00 16.15 6.54 0.11 0.32 96.23
ours 1.77 26.72 15.73 3.60 0.03 0.08 99.60
@@
Table 2: Comparison on Booster dataset in quarter resolution and balanced set. Metrics are reported separately for all regions (All), transparent regions (Trans), and nontransparent regions (NonTrans).
Method All Trans NonTrans
EPE bad2 bad3 bad5 EPE bad2 bad3 bad5 EPE bad2 bad3 bad5
DepthAnything V2 [43] 3.16 48.83 36.98 21.22 7.19 77.98 69.02 50.91 2.91 47.25 35.09 19.25
Metric3D [44] 35.55 99.70 99.32 97.73 41.55 99.37 98.93 97.91 34.89 99.71 99.32 97.59
DepthAnything V2 metric [43] 21.55 94.28 91.37 84.21 28.42 93.04 90.72 86.78 20.94 94.25 91.22 83.84
DepthPro [3] 24.44 92.98 90.23 84.25 25.65 92.98 88.90 83.05 24.14 92.91 90.16 84.19
Marigold [17] 5.99 57.90 47.13 32.63 8.46 76.33 65.90 51.52 5.72 56.79 45.87 31.26
DepthAnything V2 metric + align 5.71 62.70 48.94 32.18 12.72 77.24 68.46 54.70 5.45 62.05 48.17 31.36
Metric3D + align 3.09 43.05 29.65 16.85 8.72 76.87 64.68 47.62 2.76 41.28 27.91 15.22
DepthPro + align 4.02 53.76 40.30 23.81 6.02 64.25 55.19 40.39 3.96 53.25 39.61 23.12
Dust3R [38] 3.70 48.57 34.16 19.53 8.69 73.40 64.03 50.18 3.34 47.21 32.56 17.93
VGGT [36] 3.70 34.05 23.44 14.58 10.78 72.22 65.12 55.83 3.32 32.27 21.34 12.24
RAFT-Stereo [26] 4.08 17.61 14.87 12.17 9.55 67.84 59.43 47.46 3.13 13.10 10.70 8.63
Selective-RAFT [39] 4.05 19.48 16.64 13.57 10.08 70.02 61.79 49.64 2.92 14.94 12.38 9.93
Selective-IGEV [39] 4.52 19.23 16.51 13.84 9.22 67.00 58.99 47.21 3.52 14.69 12.28 10.20
MochaStereo [5] 3.79 16.77 14.24 11.77 9.18 66.64 58.10 45.78 2.82 12.25 10.11 8.30
StereoAnything [13] 4.36 24.13 19.20 14.50 10.54 73.48 63.53 49.37 3.29 20.09 15.37 11.08
ours 2.43 13.84 9.98 6.91 7.32 56.77 47.83 36.45 1.76 10.06 6.54 4.08
5.1 3D-Visual-Illusion Dataset
Tables 1 report a comprehensive comparison on the 3D-Visual-Illusion dataset. In this dataset, we find monocular depth estimation methods struggle significantly in inpainting, replay, and picture illusions, resulting in bad error rate in both disparity and depth spaces. Even using GT to align the monocular relative depth with absolute depth, the performance is till far bad than ours. Compared to recent foundation model (Dust3R [38] and VGGT [36]), our method shows clear advantages. Although these models combine monocular priors with explicit matching constraints, they exhibit a strong monocular bias on illusion-rich scenes, posing risks for downstream tasks like robotics and AR. Compared to stereo matching methods relying on explicit matching, our method achieves comparable results, showing that our method keep the matching constrains. We also provide visualization in Figures 5. Our method shows superior robustness to 3D visual illusions.
5.2 Booster Datset
Booster dataset [30] contains many objects with specular and transparent surfaces. We use this dataset to evalu1ate the genralization performance in mirror illusions. As shown in Table 2, our method achieves the best performance across all regions, including transparent and non-transparent areas. This demonstrates the effectiveness of our method in handling complex visual illusions, particularly in challenging scenarios involving reflections and refractions. We also provide visualization in Figure 6. Our method shows superior robustness to mirror illusions.
Figure 5: Visualization on 3D-Visual-Illusion dataset.
Figure 6: Visualization on Booster dataset
@@ 5.2 Booster Datset
Table 3: Abalation study on the Booster dataset. MF: Monocular Feature, PF: Post Fusion, APF: Adaptive Post Fusion, SF: Stereo Fusion, VLM: Vision-Language Model.
MF PF APF SF VLM epe bad2 bad3 bad4 bad5 bad6 bad7
@@ 5.3 Ablation Study
We perform ablation studies on the Booster dataset to evaluate the contribution of each module. As shown in Table 3, the baseline stereo-only model performs poorly, confirming that binocular cues alone are insufficient in challenging material regions. Adding monocular depth via simple post fusion (PF) significantly improves generalization. Incorporating high-level monocular features (MF + PF) offers further gains, highlighting the benefits of feature-level fusion. Dual GRU fusion (APF) brings marginal improvement and, in some cases, introduces noise due to redundant or conflicting updates. Using static monocular priors repeatedly in stereo GRU (SF) worsens performance, showing that naive reuse of uncertain priors can be harmful. Finally, our full VLM-based fusion yields the best results, improving accuracy by over 10 points in bad2, demonstrating its strong reasoning ability and robustness in visually ambiguous regions.

### v4

@@ 4.1 Dual-Branch Prediction Network
The dual-branch prediction network comprises a binocular disparity estimation branch and a monocular depth estimation branch. The stereo branch adopts an iterative optimization framework, extracting features from rectified image pairs and constructing a cost volume via dot product. Starting from an initial disparity of zero, a GRU-based module iteratively refines the disparity map. The monocular branch takes the left image as input and uses the frozen DepthAnything V2 [46] to predict affine-invariant inverse depth, which is disparity under unknown camera parameters. We also extract features from frozen DepthAnything V2, followed by learnable convolutions. These adapted features serve as left-view context to guide disparity refinement in the stereo branch.
@@ 4.2 VLM-Based Fusion Network
VLM Prediction Stage  In this stage, we utilize the pretrained QwenVL2-7B model [40, 1]. The visual prompt comprises the left image, binocular disparity map, and monocular disparity map. Since the textures that mislead monocular depth estimation are often too complex and diverse to be explicitly described, the language prompt is instead designed from the perspective of materials that typically confuse stereo matching. By leveraging the general reasoning capabilities of the vision-language model, we extract embedding features that help assess the relative reliability of monocular and binocular depth cues under language guidance.
@@
Specifically, the language prompts are first mapped into the embedding space. The image and text embeddings are concatenated to form $c_{e}$, which is combined with the intermediate state $y_{t_{k}}$ and passed through a stack of Transformers. Cross-attention is used to inject conditional information, and the velocity field $v_{t_{k}}$ is predicted. After multiple iterations via Equation 8, the final state $y_{1}$ is reshaped into a 2D format and decoded into image space using a variational autoencoder. The resulting feature is then concatenated with the cost volume and passed through convolution layers to predict the confidence map $I_{c}$:
@@
where $\sigma$ is the sigmoid function, $\mathcal{F}_{c}$ denotes convolution, $\mathcal{G}_{c}$ is the VAE decoder, $y^{\prime}_{1}$ is the reshaped $y_{1}$, and $V(D_{s})$ is the cost volume sampled around binocular disparity $D_{s}$.
Global Fusion Stage  Global fusion first aligns monocular disparity $D_{M}$ to the absolute/metric disparity space and then fuses the aligned monocular disparity $\tilde{D}_{m}$ with binocular disparity $D_{s}$. The alignment is achieved by affine transformation parameters $s_{m},t_{m}$:
@@
$s_{m}$ and $t_{m}$ are learned through convolutions on the concatenation of the monocular disparity $D_{m}$ and the binocular disparity $D_{s}$. Since $s_{m}$ and $t_{m}$ on low-confident regions are unreliable, we further refine the parameters by pooling $s_{m}$ and $t_{m}$ of the high-confident neighbors. After acquiring refined parameters $s_{m}$ and $t_{m}$, we compute the aligned monocular disparity $\tilde{D}_{m}$ using Equation 10. $\tilde{D}_{m}$, $D_{s}$, and $I_{c}$ are then concatenated and passed through convolutions and upsampling layers to generate the final high-resolution disparity map.
@@ 5 Experiment
We first pre-train the models on the SceneFlow dataset [28], and then fine-tune them on the virtual 3D-Visual-Illusion data. The fine-tuned models are evaluated on both the real-world 3D-Visual-Illusion data and the Booster training set [32]. We compare our method with monocular depth estimation approaches (DepthAnything V2 [46], Marigold [17], Metric3D [47], and DepthPro [3]), multi-view foundation models (Dust3R [41] and VGGT [39]), and stereo matching methods (RAFT-Stereo [27], Selective-IGEV [42], and MochaStereo [5]). For additional details on training procedures, loss functions, evaluation metrics, and prompt designs, please refer to the supplementary material.
@@
Table 1: Evaluation results on illusion regions for real-world data of the 3D-Visual-Illusion dataset. alignalign means alignment using globally shared affine parameters computed from ground truth.
Method Finetune Disparity Space Depth Space
EPE $\downarrow$ bad2 $\downarrow$ bad3 $\downarrow$ bad5 $\downarrow$ AbsRel $\downarrow$ RMSE $\downarrow$ $\delta_{1}$ $\uparrow$
DA V2 [46] $\times$ 5.81 61.45 43.18 30.57 0.14 0.15 92.86
Metric3D [47] $\times$ 12.46 94.11 91.14 82.05 0.34 0.29 48.97
DA V2 metric [46] $\times$ 16.24 92.53 87.43 75.25 0.52 0.39 48.75
DepthPro [3] $\times$ 12.26 87.08 80.60 62.43 0.28 0.25 65.92
Marigold [17] $\times$ 21.16 65.67 59.67 53.19 0.45 0.37 63.65
DA V2 metric [46] + align $\times$ 5.23 56.82 45.50 28.89 0.17 0.15 93.70
Metric3D [47] + align $\times$ 5.70 66.26 50.92 40.43 0.17 0.17 94.80
DepthPro [3] + align $\times$ 4.36 44.98 34.98 24.70 0.09 0.10 93.83
Dust3R [41] $\times$ 6.74 52.89 45.31 36.61 0.25 0.22 87.09
VGGT [39] $\times$ 6.16 53.32 44.89 37.20 0.13 0.12 78.46
RAFT-Stereo [27] $\times$ 1.62 24.32 13.20 2.97 0.04 0.06 99.18
Selective-RAFT [42] $\times$ 1.58 23.46 12.65 2.57 0.03 0.07 99.60
Selective-IGEV [42] $\times$ 1.67 24.06 13.11 2.99 0.04 0.10 99.26
MochaStereo [5] $\times$ 1.75 25.49 14.11 3.54 0.04 0.11 98.76
StereoAnything [13] $\times$ 2.41 29.00 16.15 6.54 0.11 0.32 96.23
ours $\checkmark$ 1.77 26.72 15.73 3.60 0.03 0.08 99.60
@@
Table 2: Zero-shot generalization on the balanced set of Booster dataset with quarter resolution. AllAll: all regions, TransTrans: transparent regions, NonTransNonTrans: nontransparent regions.
Method Finetune All Trans NonTrans
EPE $\downarrow$ bad2 $\downarrow$ bad3 $\downarrow$ bad5 $\downarrow$ EPE $\downarrow$ bad2 $\downarrow$ bad3 $\downarrow$ bad5 $\downarrow$ EPE $\downarrow$ bad2 $\downarrow$ bad3 $\downarrow$ bad5 $\downarrow$
DA V2 [46] $\times$ 3.16 48.83 36.98 21.22 7.19 77.98 69.02 50.91 2.91 47.25 35.09 19.25
Metric3D [47] $\times$ 35.55 99.70 99.32 97.73 41.55 99.37 98.93 97.91 34.89 99.71 99.32 97.59
DA V2 metric [46] $\times$ 21.55 94.28 91.37 84.21 28.42 93.04 90.72 86.78 20.94 94.25 91.22 83.84
DepthPro [3] $\times$ 24.44 92.98 90.23 84.25 25.65 92.98 88.90 83.05 24.14 92.91 90.16 84.19
Marigold [17] $\times$ 5.99 57.90 47.13 32.63 8.46 76.33 65.90 51.52 5.72 56.79 45.87 31.26
DA V2 metric + align $\times$ 5.71 62.70 48.94 32.18 12.72 77.24 68.46 54.70 5.45 62.05 48.17 31.36
Metric3D + align $\times$ 3.09 43.05 29.65 16.85 8.72 76.87 64.68 47.62 2.76 41.28 27.91 15.22
DepthPro + align $\times$ 4.02 53.76 40.30 23.81 6.02 64.25 55.19 40.39 3.96 53.25 39.61 23.12
Dust3R [41] $\times$ 3.70 48.57 34.16 19.53 8.69 73.40 64.03 50.18 3.34 47.21 32.56 17.93
VGGT [39] $\times$ 3.70 34.05 23.44 14.58 10.78 72.22 65.12 55.83 3.32 32.27 21.34 12.24
RAFT-Stereo [27] $\times$ 4.08 17.61 14.87 12.17 9.55 67.84 59.43 47.46 3.13 13.10 10.70 8.63
Selective-RAFT [42] $\times$ 4.05 19.48 16.64 13.57 10.08 70.02 61.79 49.64 2.92 14.94 12.38 9.93
Selective-IGEV [42] $\times$ 4.52 19.23 16.51 13.84 9.22 67.00 58.99 47.21 3.52 14.69 12.28 10.20
MochaStereo [5] $\times$ 3.79 16.77 14.24 11.77 9.18 66.64 58.10 45.78 2.82 12.25 10.11 8.30
StereoAnything [13] $\times$ 4.36 24.13 19.20 14.50 10.54 73.48 63.53 49.37 3.29 20.09 15.37 11.08
ours $\checkmark$ 2.43 13.84 9.98 6.91 7.32 56.77 47.83 36.45 1.76 10.06 6.54 4.08
5.1 Influence of 3D Visual Illusions
Table 1 presents a comprehensive comparison of the real-world data in the 3D-Visual-Illusion dataset. The real-world data is mainly composed of inpainting, picture, and replay illusions. The results of all compared methods are obtained from official code and weights, where the stereo methods use model weights pretrained on the SceneFlow dataset.
Table 3: Results of Marigold with/without finetuning on 3D-Visual-Illusion dataset.
[…]
Illusion $\times$ 21.16 65.67 59.67 0.45 63.65
Illusion $\checkmark$ 13.67 74.82 55.20 0.28 71.04
Non-illusion $\times$ 7.61 49.18 39.56 0.18 79.76
Non-illusion $\checkmark$ 7.10 55.63 44.09 0.16 77.44
Table 4: Results of stereo methods on Booster dataset with/without finetuning on 3D-Visual-Illusion data.
Method Finetune Trans NonTrans
[…]
RAFT-Stereo [27] $\times$ 9.55 67.84 59.43 3.23 13.13 10.75
RAFT-Stereo [27] SparseSparse 15.36 80.34 72.34 7.12 27.47 24.01
RAFT-Stereo [27] DenseDense 9.24 74.10 60.67 17.39 22.48 18.96
Selective-IGEV [42] $\times$ 9.50 66.85 58.90 3.60 14.74 12.34
Selective-IGEV [42] SparseSparse 9.42 64.06 54.21 5.97 14.63 12.46
Selective-IGEV [42] DenseDense 10.39 69.65 59.40 5.32 19.00 15.64
In Table 1, we can observe that monocular depth estimation methods struggle notably with inpainting, replay, and picture illusions, leading to large errors in both disparity and depth spaces. Even when ground-truth alignment is used to convert relative depth to absolute scale, their performance remains significantly worse than stereo methods and ours. As for recent foundation models, such as Dust3R [41] and VGGT [39], they exhibit a strong monocular bias in illusion-rich scenes. In comparison with stereo matching methods depending on explicit correspondence, our method achieves comparable results, indicating that it preserves strong matching constraints. Qualitative results in Figure 6 further illustrate the serious influence of 3d visual illusions on SOTA depth estimation methods. More visualization, please refer to our supplemental materials.
The Booster dataset [32] includes many objects with specular and transparent surfaces, making it well-suited for evaluating generalization to mirror illusions. As shown in Table 2, monocular methods achieve the best performance, especially DepthPro. In contrast, binocular methods are easily fooled by mirrors and specular objects. Qualitative results in Figure 6 further show that mirror illusions pose a serious challenge to SOTA binocular methods. See supplemental materials for more visualizations.
Here, we also compare classical binocular and monocular methods with/without finetuning on 3D-Visual-Illusion data. Table 4 shows Marigold’s performance on illusion and non-illusion regions of real-world 3D-Visual-Illusion data. We omit DepthAnything V2 results, as its official code fails to converge on the virtual data. This is because its official implementation trains on metric depth, which is sensitive to scale variations, while our virtual data varies significantly in scale. In contrast, the official code of Marigold supports training on affine-invariant depth, ensuring stable learning. Table 4 shows that fine-tuning improves most metrics on illusion regions, suggesting Marigold adjusts its predictions toward planar surfaces, but at the cost of degraded performance on non-illusion regions. This supports our hypothesis that monocular depth models rely on fixed mappings from texture cues (e.g., shape, shadow, perspective, defocus), making it difficult to distinguish illusion textures from real object textures. The worsened bad2 metric in illusion regions indicates incomplete overfitting, while the limited EPE gain in non-illusion regions likely stems from correcting a few extreme outliers rather than consistent improvement.
Table 4 presents the performance of SOTA stereo models with and without finetuning on the 3D-Visual-Illusion dataset. The $Sparse$ and $Dense$ represent different augmentation strategies used during training. The results show that all stereo models achieve only limited improvement on transparent regions after finetuning, while suffering a significant performance drop on non-transparent regions. This indicates that standard stereo architectures cannot effectively learn from such data. When finetuning on our virtual illusion data, different illusion types are inherently conflicting: mirror illusions rely on spatial context (i.e., monocular priors) for accurate depth estimation, whereas inpainting, picture, replay, and holography illusions deliberately mislead models by distorting these priors. Thus, features learned from mirror illusions are compromised when the model is trained on other illusion types, leading to conflicting learning of monocular priors. Moreover, since we assume a flat plane to rectify disparity during the generation of virtual illusion data, the finetuned stereo models tend to produce overly flat disparities. This results in a slight improvement on transparent glass regions but severe degradation on other non-transparent and non-planar objects.
In addition to illusion scenes, we also present the performance of our model in the Middlebury dataset. We compare our model with several SOTA stereo-based approaches using metric disparity space over the entire image, without restricting the maximum disparity range. Table 5 shows that our model does not degrade in performance on these mundane scenes. On the contrary, it achieves improvements, particularly in terms of EPE.
Table 5: Zero-shot generalization on Middlebury dataset at half resolution. Evaluation is conducted in metric disparity space over the entire image, without restricting maximum disparity.
Metric Selective-RAFT Selective-IGEV MochaStereo StereoAnything RAFT-Stereo Ours
EPE $\downarrow$ 2.34 2.59 2.66 2.89 1.92 1.50
Bad-2 $\downarrow$ 12.04 11.79 10.18 11.93 12.60 11.79
Figure 5: Visualization on our dataset.
Figure 6: Visualization on Booster dataset.
@@ 5.1 Influence of 3D Visual Illusions
Table 6: Ablation study on the Booster dataset. MF: Monocular Feature, PF: Post Fusion, APF: Adaptive Post Fusion, SF: Stereo Fusion, VLM: Vision-Language Model.
MF PF APF SF VLM EPE $\downarrow$ bad2 $\downarrow$ bad3 $\downarrow$ bad4 $\downarrow$ bad5 $\downarrow$ bad6 $\downarrow$ bad7 $\downarrow$
@@ 5.2 Ablation Study
We conduct ablation studies on the Booster dataset to evaluate the contribution of each module and various fusion strategies. As shown in Table 6, the baseline stereo-only model performs poorly, indicating that binocular cues alone are insufficient in regions with challenging materials. Introducing monocular depth through simple post fusion (PF, where fusion is guided by confidence generated from image features) significantly improves generalization. Incorporating monocular features into the stereo branch (MF + PF) further improves performance on the bad metrics, although it slightly degrades the EPE error rate, which indicates better overall geometry but more severe outlier shifts. Adaptive post fusion (APF) employs two independent GRUs to iteratively update the monocular and binocular disparity during fusion, where each other’s disparity is used as update guidance. Although this strategy can bring some improvements, it may introduce noise due to inconsistent updates between the two branches. SF uses a single GRU to fuse binocular disparity with fixed monocular disparity, which makes performance worse, highlighting the risks of naïvely reusing uncertain priors. Finally, our full VLM-based fusion approach achieves the best performance, improving the bad2 metric by over 10 points. This demonstrates the strong reasoning capability of the VLM in handling visually ambiguous regions and its effectiveness in guiding reliable depth fusion. We also evaluate the VLM’s confidence by comparing the predicted confidence maps with the disparity error maps, as defined in Equation 14. This analysis is conducted on the Booster dataset under a zero-shot generalization setting. The results show that the error rate of our confidence estimation is approximately $20\%$, demonstrating a strong generalization ability, even in previously unseen illusory scenes.
5.3 Discussion
Illusion Effect on Different Depth Paradigms: (1) Monocular estimation relies on texture-based cues (e.g., shape, perspective, shadow, defocus) learned from RGB image. When these cues are artificially simulated on flat surfaces (inpainting, pictures, replays, holograms), the model is easily misled, producing incorrect depths. Mirror illusions, however, can often be resolved through scene-level context learned from large-scale data. (2) Stereo estimation instead depends on pixel-wise correspondence. In mirror scenes, reflections overlap with real surfaces, causing ambiguous matches and depth errors. For inpainting, pictures, replays, and holography, stereo matching remains effective due to cross-view texture consistency.
Limitations and Future Work: (1) The virtual data generation pipeline relies on manual semantic segmentation, which is labor-intensive and time-consuming. Given the lack of 3D visual illusion data, and the fact that existing detectors, segmentation models, and VLMs are often fooled, even humans will be fooled in complex cases, manual collection remains a practical step at this stage. Developing an automatic pipeline is an important future direction. (2) The real-world subset currently covers only a limited range of illusions (inpainting, picture, and replay). We plan to extend it to broader types and more diverse real-world scenes. (3) The VLM-driven fusion is effective but computationally expensive. Designing lighter, more efficient fusion methods is worth further exploration. (4) Our study focuses on pure illusions without compositing. Future challenges include combinations of multiple illusions, semantically ambiguous objects, and entirely novel illusion types.